OpenStudy (anonymous):

Log_2 * 3.75 = x (solve)

OpenStudy (unklerhaukus):

do you mean \[log_2(3.75)=x\]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

\[\frac{\ln(3.75)}{\ln(2)}\] and a calculator